Circular No. 9001

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 X1 (HILL) R. E. Hill reports his discovery of a comet on Catalina Sky Survey CCD

images obtained with the 0.68-m Schmidt reflector (discovery observation tabulated below), the comet noted from co-added exposures as has having a nuclear condensation of size 8′′–10′′ with a diffuse coma of diameter 30′′ and a broad, diffuse tail 15′′-20′′ long in p.a. 290^{o}–300^{o}. Four additional co-added 60-s exposures by Hill on Dec. 5.2 UT in 2′′-4′′ seeing show a similar appearance, with the nuclear condensation measured as 10′′–12′′ across. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD observers have noted the object’s cometary appearance. R. Ligustri (Talmassons, Udine, Italy), observing remotely with a 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., finds a coma diameter of ∼ 25′′ on twelve co-added 180-s exposures taken on Dec. 5.3; using the same instrumentation, E. Guido, G. Sostero, and P. Camilleri report that 24 co-added 120-s unfiltered exposures from Dec. 5.4 show a compact inner coma ∼ 10′′ in diameter, and an extremely faint, fan-shaped, broad tail ∼ 30′′ long toward the northwest. R. A. Kowalski (Mt. Lemmon 1.5-m reflector, Dec. 5.37) reports a coma ≈ 12′′ in diameter and a very faint, low-contrast, fan-shaped tail in p.a. 290^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
Dec. 4.34227  5^{h}24^{m}20.^{s}26  +3^{o}17′52′′.2  17.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-X42.

T = 2009 May 20.504 TT  ω = 44.^{o}947
Ω = 108.322 }2000.0
q = 1.61769 AU  i = 23.774

COMET C/2008 O7 (SOHO) Further to IAUC 8987, below are tabulated the discovery data for an

additional near-sun presumed comet found on SOHO website images. Not a member of any known dynamical group, C/2008 O7 was very diffuse and perhaps slightly elongated (mag ∼ 7.5–8.0); based on its appearance, K. Battams remarks that it is unlikely to have survived.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 O7 July 18.921 7^{h}54.^{m}6 +19^{o}12′ C2 BZ 2008-T10
2008 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green